OpenStudy (anonymous):

Find the volume of the solid formed by rotating the region bounded by: y=x^2+2 and y=x+8 around the line, x=4. Leave the answer in terms of pi. (calculus)

OpenStudy (anonymous):

use formula for the volumes by cylindrical shells... 1) Find point in intersection of two graphs - will give you int limits: x^2+2=x+8 I got x=-2 & x= 3 2) v= int (from -2 to 3) 2*pi * (4-x)*(x+8 - x^2-2)dx = ... should be OK from this point let me know if there is a problems with solution